OpenStudy (anonymous):

Find the indicated root. Leave the answer as radicals. (Do not change to decimals!) 3square root 81 (3 is the index number)

OpenStudy (anonymous):

\[\sqrt[3]{81}=\left( 81 \right)^{1/3}=(9^{2})^{1/3}\]

OpenStudy (anonymous):

can you do it from here?

OpenStudy (anonymous):

yeah thank you!